OpenStudy (nwolf):

Write an equation of the given point and perpendicular to the given line (9,-7) 8x+7y=9 the equation of the line is y=

OpenStudy (anonymous):

8x +7y = 9 7y = -8x +9 y = -8/7x + 9/7 m = -8/7 perpendicular slope is 7/8 y + 7 = 7/8 (x - 9) y = 7/8x -63/8 -56/8 y = 7/8x - 119/8